OpenStudy (anonymous):

Can someone please help me simplify 5root16 - 3root4? Thanks!

OpenStudy (anonymous):

sq root??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(\Large 5\sqrt{16} -3\sqrt{4}=5\sqrt{4^2}-3\sqrt{2^2}=5\times4-3\times2=14\)

OpenStudy (anonymous):

Thank you, so would that be just 14 for a final answer? Wouldn't we incorporate the exponents somehow in the final answer? if we just do 5*4, what is happening to the fact that the 4 is really a 4^2??

OpenStudy (anonymous):

the final answer: \(\Large 5\sqrt{16} -3\sqrt{4}=5\times4 - 3\times 2=14\)

OpenStudy (anonymous):

So do the exponents (^2) cancel each other out because they match or what happens with them?

OpenStudy (anonymous):

because it is a easy calculation, so we just do that

OpenStudy (anonymous):

oh, ok. Thank you for all the help!